OpenStudy (anonymous):

Input in standard form the equation of the given line. The line that passes through (-1, -3) and (2, 1)

OpenStudy (anonymous):

IMPORTANT LINE RELATED EQUATIONS TO KNOW AND MEMORIZE slope formula m= slope/ gradiant -- same thing \[m=\frac{y_2-y_1}{x_2-x_1}\] standard formula \[Ax+By=C\] point-slope formula \[y-y_1=m(x-x_1)\] slope-intercept formula b= y-intercept -- in the form of (0,y) \[y=mx+b\]

OpenStudy (anonymous):

i tried that but the answer doesnt seem to be correct when i plug the numbers back in

OpenStudy (anonymous):

what is the slope?

OpenStudy (anonymous):

my slope was 4/3

OpenStudy (anonymous):

good, now take the slope and one of the points and substitute it into the point slope formula

OpenStudy (anonymous):

this is what i got y+3=4/3(x+1)

OpenStudy (anonymous):

ok good now convert it into standard form

OpenStudy (anonymous):

first start by distributing the 4/3

OpenStudy (anonymous):

i got y+3=4/3x+4/3

OpenStudy (anonymous):

good now change it standard form

OpenStudy (anonymous):

ok i got -4/3x+y=5/3 the only problem i have with that answer is that for the website i have to use for my summer work is that there are boxes in which i have to input the numbers by dragging them in the boxes and there is no box to input a fraction

OpenStudy (anonymous):

ok, well that means we need to get rid of the fraction this can be done if you find the greatest common denominator of the equation and multiplying it to both sides in other words multiply the entire equation ( both sides) by 3

OpenStudy (anonymous):

oh wait you made a mistake it should be -5/3 not 5/3

OpenStudy (anonymous):

oh yes i forgot to put that

OpenStudy (anonymous):

ok so i got -4x+3y=-5

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you very much :)

OpenStudy (anonymous):

no problem it looks like you already know how to solve problems like these, so just take it step by step and slowly next time

OpenStudy (anonymous):

yeah this is supposed to be algebra 1 review to get into trig/precalc

OpenStudy (anonymous):

well, if you have any more questions we'll be happy to help unless they're problems from exams

OpenStudy (anonymous):

oh no these are just homework help questions. thank you again